What is the thermal efficiency of an air standard gas turbine in %, if the density of the gas increases 345 % during compression and the heat capacity ratio is 1.39?
1
Expert's answer
2020-05-20T09:27:48-0400

The thermal efficiency of an air standard gas turbine in % is


"\eta_\text{th}=\bigg(1-\frac{1}{\lambda^{\gamma-1}}\bigg)\cdot100\%,\\
\space\\
\eta_\text{th}=\bigg(1-\frac{1}{(345/100)^{1.39-1}}\bigg)\cdot100\%=38.3\%."